OpenStudy (anonymous):

int[xsinx(cosx)^4]dx from 0 to pi/4. Is this tabular integration? I'm not sure how to do that if so.. thank you.

OpenStudy (anonymous):

Could yo rewrite the problem? BTW, Do you mean definite integration or integration by parts?

OpenStudy (anonymous):

sorry, i wrote the problem incorrectly.

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi/4}(\theta*\sin(\phi)\cos^4(\phi))d \phi\]

OpenStudy (anonymous):

where theta is treated as a constant

OpenStudy (anonymous):

okay, so do you know integration by parts?

OpenStudy (anonymous):

If this case, u would be sin(tetha) and dv would be cos^4(tetha)

OpenStudy (anonymous):

\[\int\limits_{}^{}(udv)=uv-\int\limits_{}^{}v du\]

OpenStudy (anonymous):

ok, would i have to do integration by parts multiple times bc of the power to 4?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

oooh